Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night Report Page - 20120922

Observers : Iair Arcavi, Yen-Chen Pan. 
Data reduction team : K. Maguire, M. Sullivan, A. Gal-Yam, O. Yaron

Some clouds during the second half of the night. Not photometric (first half might be ok though).
SOFI photometric standard: bd174708003
EFOSC spectroscopic standards: Feige110, gm71
SN2009ip [SOFI: JHK Bg Rg]
Seeing ~0.9". Hip110512 taken before the Bg and after the Rg spectra.
Note that we stopped at one point to make sure the object was in the slit (since the trace seemed faint) and then continued.
The object is isolated and clearly seen to enter the slit. We're using the deepest OB under good conditions. 
Note the total time to complete the SOFI OB for this target is ~6 hours.
SN2012dy [SOFI: JHK]
Seeing ~0.8".
SN2012ec [SOFI: JHK Bg Rg]
Cancelled due to noise problem (again) - same quadrant as last time (top right). Changing the rotator angle didn't work. 
Next fix is going to the telescope and readjusting some cards. This can take 30-40 minutes so we switched to EFOSC and 
will ask the day crew to take care of it for the next nights.
SOFI --> EFOSC
LSQ12dlf [EFOSC: BVRi gm13]
Seeing ~0.9".
SN2012ec [EFOSC: gm11 gm16]
Seeing ~1".
Comments